There is a meteor shower. One meteor appears at A(7,8) and disappears at B (1,4).
Calculate the gradient of the meteor's flight path.
calculate the acute angle between the flight path and the horizon (OX) correct to the nearest degree.
A second meteor, on a parallel path, starts at C(6,2) and disappears below the horizon at D. Find the co-ordinates of D.

Answer by josmiceli(9817)

slope of path between (7,8) and (1,4) = (8-4) / (7-1) = 2/3
angle between path and horizon = inverse tan(2/3) = 33.7 degrees
2nd meteor path through (6,2)
y = (2/3)x + b
2 = (2/3)6 + b
2 = 4 + b
b = -2
y = (2/3)x - 2
at horizon point DE is (x,0), find x coordinate
0 = (2/3)x - 2
(2/3)x = 2
x = 3
at horizon point DE is (3,0)